OpenStudy (anonymous):

for f(x) = 1/x^2-3, find f (3) f (2+h)

OpenStudy (anonymous):

For f(3), plug in 3 where the x is.

OpenStudy (anonymous):

For f(2+h) your answer will have 1 divided by a quadratic of h in it.

OpenStudy (anonymous):

so 1 is 9 and idk 2

OpenStudy (anonymous):

2+h just replaces x. It's basically a change in variable

OpenStudy (anonymous):

so f = 2+h

OpenStudy (anonymous):

is should be in the form\[f(2+h)=\frac{ 1 }{ (2+h)^2 }-3\]

OpenStudy (anonymous):

since x is replaced by 2+h

OpenStudy (anonymous):

you understand how i did this, right?

OpenStudy (anonymous):

oh okay, what about the first one? 9 is correct?

OpenStudy (anonymous):

No you have to solve for \[f(3)=\frac{ 1 }{ (3)^2 }-3\]

OpenStudy (anonymous):

and whats that? :/

hartnn (hartnn):

is your f(x) \(=\dfrac {1}{x^2-3}\) ??

OpenStudy (anonymous):

\[f(3)=\frac{ 1 }{ (3)^2 }-3=\frac{ 1 }{ 9 }-3=-\frac{ 26 }{ 9 }\]

OpenStudy (anonymous):

so that's the answer? :S

OpenStudy (anonymous):

I'm really not good at math sorry!! :/

OpenStudy (anonymous):

lol it's really about trying and grasping the topics you learn in school and continuously practicing them. No one is born knowing how to completely do math. Everything is practice and applying what you have learned But that is the answer